OpenStudy (rainbow_dash):

one more question on the same subject. A student guesses the answers to 6 questions on a true-false quiz. Find the probability that the indicated number of guesses are correct: no more than 2 (Hint: No more than 2 means exactly 0 or exactly 1 or exactly 2.)

OpenStudy (anonymous):

then your job is to compute exactly 0 right, exactly 1 right, and subtract those from one

OpenStudy (anonymous):

of course it is math, that is why she is guessing

OpenStudy (unklerhaukus):

i see now

OpenStudy (anonymous):

exactly none means all wrong only one way to get all wrong, namely wrong wrong wrong wrong wrong wrong probability is therefore \[\frac{1}{2^6}\] for all wrong

OpenStudy (rainbow_dash):

would it be 2/2^6 for 1 right?

OpenStudy (anonymous):

one wrong there are 6 possibilities

OpenStudy (anonymous):

or one right, same thing

OpenStudy (anonymous):

R W W W W W W R W W W W W W R W W W W W W R W W W W W W R W W W W W W R

OpenStudy (rainbow_dash):

i think i got my answer, total is gonna be .08-.09

OpenStudy (anonymous):

otherwise known as \(\dbinom{6}{1}=6\)

OpenStudy (anonymous):

so it is \[1-\frac{1}{2^6}-\frac{6}{2^6}=1-\frac{7}{2^6}\]

OpenStudy (rainbow_dash):

Excellent, thanks :)